OpenStudy (anonymous):

A plane accelerates from rest at a constant rate of 5.00 m/s^2 along a runway that is 1800 m long. Assume that the plane reaches the required takeoff velocity at the end of the runway. What is the time needed to take off?

OpenStudy (anonymous):

i got 134.16 sec??

OpenStudy (anonymous):

oops never mind :)

OpenStudy (anonymous):

I think you take 1800/5 to see how many 5m are in the 1800m which is 360 so it would take 360 (5m) to get to 1800 so s^2 = 360 square root both sides about 18.97 seconds... Not sure though

OpenStudy (anonymous):

s = ut + (1/2)at^2 => 1800 = 0t + (5/2) t^2 => t^2 = 3600/5 =>t = sqrt(3600/5) = sqrt(720) = 12sqrt(5)

OpenStudy (anonymous):

nice thanks Guys!